Find an equation of the line containing (-1, 3) and perpendicular to the line containing (3, -5) and (-2, 7).
Answer by Alan3354(69443)

Find an equation of the line containing (-1, 3) and perpendicular to the line containing (3, -5) and (-2, 7).
---------------------------
This is a 2 step process. 1st find the slope m1, of the line. Then find the line perpendicular. Perpendicular means the slope is the negative inverse of the slope m1.
----------
m1 = (y2-y1)/(x2-x1)
m1 = (-5-7)/(3 -(-2))
m1 = -12/5
So the slope of the perpendicular is +5/12
Use the slope-intercept eqn
y-y3 = m*(x-x3) where (x3,y3) is the point (-1,3)
y-3 - 5/12(x-(-1)
y-3 = 5/12x + 5/12
y = 5x/12 + 3 5/12 (slope-intercept form)
or y = (5/12)x +41/12
-----------
12y = 5x + 41
5x - 12y = -41 (standard form)